Frequency of tuning fork is 500 hz and the velocity of sound in air is 300 m/s the distance travelled by a sound while fork  execute 200 vibration

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Wavelength = 300/500 = 3/5
 
Hence
 
Distance travelled = 3/5 * 200 = 120 m
 
Hope it helps
 
 
Regards
Arun